Sweet Corporation’s charter authorized issuance of 105,000 shares of $10 par value common stock and 51,000 shares of $50 preferred stock. The following transactions involving the issuance of shares of stock were completed. Each transaction is independent of the others.

1. Issued a $9,500, 9% bond payable at par and gave as a bonus one share of preferred stock, which at that time was selling for $117 a share.
2. Issued 480 shares of common stock for equipment. The equipment had been appraised at $7,300; the seller’s book value was $5,600. The most recent market price of the common stock is $15 a share.
3. Issued 376 shares of common and 98 shares of preferred for a lump sum amounting to $10,100. The common had been selling at $13 and the preferred at $60.
4. Issued 180 shares of common and 53 shares of preferred for equipment. The common had a fair value of $15 per share; the equipment has a fair value of $6,700.


Record the transactions listed above in journal entry form. (Round Round intermediate calculations to 6 decimal places, e.g. 0.546872 and final answers to 0 decimal places, e.g. $38,487. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

1.
2.
3.
4.
Answer #1

Journal Entries Sl.No 1 (see workings 1) Credit $ $ Debit 9,500 117 Accounts Titles & Explanation Cash Discount on bonds payaWorkings 1 Preferred stock = No. of preference shares issued * issue price = 1 * 50 = $50 Paid in capital in excess of par prWorkings 3 Common stock = No. of common stocks issued * issue price = 376 * 10 = $ 3760 Paid in capital in excess of par commWorkings 4 Common stock = No. of common stocks issued * issue price = 180 * 10 = $ 1800 Paid in capital in excess of par comm